In the following one solution of a second y1 order linear homogene DE is given. Find the second linearly independent solution y2 using the method of reduction of order.


  1. 2x²y" + 3xy'-y=0 ,.
  2. (1-x2)y"-2xy'+2y=0 ,
  3. x2y"+2xy'-2y=0,
  4. x2y"+3xy'+y=0,
  5. x2y"-x(x+2)y'=0,
1
Expert's answer
2021-07-13T05:14:42-0400

1.

Euler Equation.

"y=x^m"


Then:

"x^2m(m-1)x^{m-2}+3\/2xmx^{m-1}-x^m\/2=0"


Characteristic Equation:

"m(m-1)+\frac{3}{2}m+\frac{1}{2}=0"

or

"m^2+\frac{1}{2}m+\frac{1}{2}=0"


"m_{1,2}=\frac{-1\pm i\sqrt{3}}{4}"


Thus, the general solution is

"y=x^{-1/4}[c_1cos(\sqrt{3}ln|x|/4)+c_2sin(\sqrt{3}ln|x|/4)]"


2.

"(1+x^2)y''-2xy'+2y=0"


"y_1=x"


"y''-2xy'\/(1+x^2)+2y\/(1+x^2)=0"


"y_2=\int\frac{e^{-\int p(x)dx}}{y^2_1}dx"


where p(x)=-2x/(1+x2)


Then:


"\int \frac{2x}{1+x^2}dx=ln|1+x^2|"


"y_2=\int \frac{dx}{x^2(1+x^2)}=-1/x+arctanx"


"\frac{1}{x^2(1+x^2)}=\frac{A}{x}+\frac{B}{x^2}+\frac{Cx+D}{1+x^2}"


"Ax(1+x^2)+B(1+x^2)+Cx^3+Dx^2=1"


"A+C=0"

"B+D=0"

"A=0,B=1,D=-1,C=0"


3.

x2y"+2xy'-2y=0

"y=x^m"

The characteristic equation:

"m(m-1)+2m-2=0"


"m_{1}=\frac{-1+ 3}{2}=1,m_2=-2"


The general solution is

"y=c_1x+c_2x^{-2}"


4.

x2y"+3xy'+y=0

"y=x^m"

The characteristic equation:

"m(m-1)+3m+1=0"


"m_{1,2}=\frac{-2}{2}=-1"


"y_1=x^{-1},y_2=x^{-1}ln|x|"


5.

x2y"-x(x+2)y'=0

"y'=z,y''=dz\/dx"

Then:

"x^2z'-x(x+2)z=0"


"\frac{dz}{z}=\frac{x(x+2)}{x^2}dx"


"ln|z|=x+2ln|x|+c_1"


"z=x^2e^{c_1x}"


"y=\intop x^2e^{c_1x}dx=\frac{e^{c_1x}(c_1^2x^2-2c_1x+2)}{c_1^3}+c_2"